Three capacitors each of capacity C are connected in series the resultant capacity wil be

Grade:11

1 Answers

Narendran
7 years ago
* C/3
1/Cseries=1/C1+1/C2+1/C3….….….….. Put C1=C2=C3=C ;
Then try youself for the required    equation!!
 
WHAT IS BEHIND.
In series connection sum of individual potential drops equals the total potential drop V applied across the input and output wires: i.e., V=V1 +V2 The equivalent capacitance of the pair of capacitors is again Ceq=Q/V. Thus,
1/Ceq=V/Q=(V1+V2)/Q=V1/Q+V2/Q
1/Ceq=1/C1+1/C2.
Here, we have made use of the fact that the charge Q is common to all three capacitors. Hence, the rule is:
The reciprocal of the equivalent capacitance of two capacitors connected in series is the sum of the reciprocals of the individual capacitances.